# Wave–Particle Complementarity as the Optimal Limit of Unambiguous Quantum-State Discrimination

### v2

@@ II ZWM Idler States and Hypothesis Testing
IDP interpretation of visibility.— The two idler states in the low-gain ZWM interferometer are $|\phi_{1}\rangle=r|1,0\rangle+t|0,1\rangle$ and $|\phi_{2}\rangle=|0,1\rangle$, with overlap $|\langle\phi_{1}|\phi_{2}\rangle|=|t|=V$. In the standard complementarity framework, visibility $V$ and particle-like distinguishability
$D=\sqrt{1-V^{2}}$
are determined solely by this overlap.
We emphasize that the quantities appearing in our analysis have distinct operational meanings. Englert distinguishability $D$ is conceptually different from the two outcome probabilities of the IDP measurement. For IDP, the optimal conclusive success probability,
$P_{S}^{\mathrm{opt}}=1-|\langle\phi_{1}|\phi_{2}\rangle|,$ (10)
and the optimal inconclusive (failure) probability,
$P_{I}^{\mathrm{opt}}=|\langle\phi_{1}|\phi_{2}\rangle|,$ (11)
are complementary outcomes satisfying $P_{S}^{\mathrm{opt}}+P_{I}^{\mathrm{opt}}=1$. Thus, $P_{S}^{\mathrm{opt}}$ and $P_{I}^{\mathrm{opt}}$ are not independent quantities. Importantly, the distinguishability $D$ arises from minimum-error discrimination and is therefore not equal to either $P_{S}^{\mathrm{opt}}$ or $P_{I}^{\mathrm{opt}}$, which are the complementary outcome probabilities of unambiguous discrimination.
For pure markers in the ZWM geometry, the visibility equals the IDP failure probability,
$V=|\langle\phi_{1}|\phi_{2}\rangle|=P_{I}^{\mathrm{opt}},$ (12)
yielding the operational reformulation of complementarity,
$D^{2}+\big(P_{I}^{\mathrm{opt}}\big)^{2}=1.$ (13)
Thus, wave–particle duality in induced coherence can be understood as a statement about the optimality of a zero-error quantum measurement on the idler mode. The explicit construction of the optimal POVM $\{\Pi_{0},\Pi_{1},\Pi_{2}\}$ implementing the IDP measurement is presented in Supplemental Material.
@@
The same reasoning applies to any balanced two-path interferometer in the single-excitation regime with pure which-path markers $|\mu_{1}\rangle$ and $|\mu_{2}\rangle$. The fringe visibility is determined by their overlap,
$V=|\langle\mu_{1}|\mu_{2}\rangle|,$
@@
$D=\sqrt{1-|\langle\mu_{1}|\mu_{2}\rangle|^{2}}.$
For equal priors, the optimal IDP strategy yields $P_{I}^{\mathrm{opt}}=|\langle\mu_{1}|\mu_{2}\rangle|$, so that
$D^{2}+\big(P_{I}^{\mathrm{opt}}\big)^{2}=1$
@@
Figure 2: (a) Operational complementarity between visibility $V$, distinguishability $D$, and the optimal IDP failure probability $P_{I}^{\mathrm{opt}}$ for pure marker states with overlap $s=|\langle\mu_{1}|\mu_{2}\rangle|$ (see Appendix D). The curves satisfy $D^{2}+V^{2}=1$ and $V=P_{I}^{\mathrm{opt}}$, identifying wave–particle duality with the optimality condition of unambiguous state discrimination. (b) Comparison between distinguishability obtained via the coincidence-based method of Torres et al. [Torres2024] (solid) and via optimal IDP discrimination on the idler (dashed). The coincidence method saturates at $D_{\mathrm{coinc}}(T=1)\approx 0.52$, whereas IDP extracts greater distinguishability for intermediate $0<T<1$. Both approaches correctly give $D=0$ at $T=1$, where the marker states become identical. The advantage of IDP therefore appears only for partially overlapping markers.
Figure 2 provides a direct operational comparison between distinguishability extracted from signal-arm visibility (Torres et al.) [Torres2024] and distinguishability obtained via optimal IDP discrimination performed on the idler. In panel (a), we plot the complementarity relation for pure which-crystal markers, showing that the visibility $\mathcal{V}$, the optimal IDP failure probability $P_{I}^{\mathrm{opt}}$, and the particle-like distinguishability $D$ are constrained by the operational identity $\mathcal{V}=P_{I}^{\mathrm{opt}}$ and $D^{2}+\mathcal{V}^{2}=1$. This demonstrates that wave–particle duality is exactly equivalent to an optimal unambiguous state discrimination task for the which-path markers.
Panel (b) compares distinguishability extracted from the coincidence-based method of Torres et al. [Torres2024] (solid curve) with the distinguishability achievable via optimal Ivanovic–Dieks–Peres (IDP) unambiguous discrimination on the idler (dashed curve), both expressed in the same metric. In the Torres approach, $D$ is inferred from the imbalance between second-order correlations $g^{(2)}_{13}$ and $g^{(2)}_{23}$, and therefore saturates at a finite value $D_{\mathrm{coinc}}(T=1)<1$ even when the idler transmissivity $T$ is unity. In contrast, the IDP strategy operates directly on the which-crystal idler marker states and achieves the optimal extraction of available which-path information. Both methods converge to $D=0$ when $T=1$, since the idler marker states become identical and no measurement—IDP or otherwise—can distinguish the emission events. The advantage of IDP emerges in the intermediate regime $0<T<1$, where coincidence-based extraction underestimates the distinguishability available in principle. This highlights that path information in induced coherence is fundamentally limited by measurement strategy rather than interferometer geometry.
Operational Interpretation of Visibility as IDP Failure Probability
The result $V=P_{I}^{\mathrm{opt}}$ gives visibility a concrete information–theoretic meaning, which we now make explicit. In the induced-coherence regime, the idler modes emitted by the two SPDC crystals prepare the nonorthogonal marker states $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$. Since these states satisfy $\langle\phi_{1}|\phi_{2}\rangle\neq 0$, quantum mechanics forbids any measurement from identifying the emitting crystal with zero error and nonzero success probability. The only strategy that guarantees error-free identification is unambiguous discrimination, in which the measurement is constrained to yield conclusive outcomes without ever misidentifying the state, at the cost of introducing a third, inconclusive outcome; see Appendix D. The POVM implementing this strategy has three elements $\{\Pi_{0},\Pi_{1},\Pi_{2}\}$, corresponding to conclusive identification of $|\phi_{1}\rangle$, conclusive identification of $|\phi_{2}\rangle$, and an inconclusive result.
For two pure states with equal priors, the optimal IDP measurement minimizes the inconclusive probability, which is bounded below by the state overlap itself:
$P_{I}^{\mathrm{opt}}=|\langle\phi_{1}|\phi_{2}\rangle|.$ (14)
This expression is a fundamental quantum limit: no physically allowed measurement can unambiguously discriminate $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$ with a smaller failure probability without violating the zero-error constraint. Equation (14) therefore identifies the state overlap as the operational quantity governing the difficulty of the discrimination task.
In the ZWM interferometer, the same overlap controls the visibility of single-photon interference in the signal arm, $V=|\langle\phi_{1}|\phi_{2}\rangle|$. Thus, the single-photon visibility coincides exactly with the minimum inconclusive probability of an error-free measurement on the idler:
$V=P_{I}^{\mathrm{opt}}.$ (15)
This equality shows that the conventional interference visibility is not merely a geometric measure of coherence but the optimal failure probability in a well-defined quantum decision problem. The wave–particle duality relation $D^{2}+V^{2}=1$ is therefore understood as a measurement optimality theorem: the distinguishability $D$ reflects the maximum achievable probability of conclusive identification, while the visibility $V$ corresponds to the unavoidable inconclusive probability of the optimal unambiguous discrimination protocol.
This operational perspective clarifies the role of coherence in induced interference and unifies two concepts that have historically been treated separately. Rather than simply quantifying fringe contrast, visibility characterizes the fundamental decision-theoretic limitation imposed by the nonorthogonality of the idler states. Consequently, wave–particle duality in the ZWM interferometer emerges as the boundary between error-free discrimination and coherent interference, established by the IDP optimality condition.
III Experimental Protocol

### v3

@@ II Induced-coherence interferometer marker states and complementarity
Previous ZWM experiments have measured visibility and which-path information as evidence of the complementarity relation $D^{2}+V^{2}=1$ [Torres2024], but they do not address the fundamental limit on identifying the emitting crystal in a single-shot measurement. This motivates the operational question: How well can one unambiguously (error-free) discriminate which crystal produced the photon pair? Because the conditional idler states are nonorthogonal, perfect identification is impossible, and any zero-error strategy must rely on unambiguous discrimination.
III Quantum decision strategies for which-crystal discrimination
In the ZWM setup, low-gain regime induced-coherence scenario, the idler subsystem encodes the binary hypothesis which crystal fired, more details see Appendix B. We denote the two hypotheses by
$\begin{split}\mathcal{H}_{A}&:\ \hat{\rho}=\hat{\rho}_{A}\ \ \text{(crystal A fired)},\qquad\text{prior }p_{0},\\ \mathcal{H}_{B}&:\ \hat{\rho}=\hat{\rho}_{B}\ \ \text{(crystal B fired)},\qquad\text{prior }p_{1},\end{split}$ (10)
with $p_{0},p_{1}\geq 0$ and $p_{0}+p_{1}=1$.
In the idler mode $I$ performs a dichotomous POVM $\{\Pi_{A},\Pi_{B}\}$ with outcomes $\mu\in\{0,1\}$ satisfying
[…]
where outcome $\Pi_{A}$ corresponds to the decision “A fired” ($\mu=0$) and outcome $\Pi_{B}$ corresponds to the decision “B fired” ($\mu=1$).
In the Conditional error probabilities, If $\mathcal{H}_{A}$ is true (state $\hat{\rho}_{A}$ was prepared), an error occurs when crystal outputs “B”, i.e., when outcome $\Pi_{B}$ clicks. Similarly, if $\mathcal{H}_{B}$ is true (state $\hat{\rho}_{B}$ was prepared), an error occurs when crystal outputs “A”, i.e., when outcome $\Pi_{A}$ clicks:
[…]
The Bayes average error probability of which crystal fired is
[…]
IV Minimum-error strategy: Helstrom bound
In the low-gain ZWM interferometer, the two hypotheses “crystal A fired” and “crystal B fired” are encoded in the conditional idler states see more details Appendix C
[…]
$|.s:=|\innerproduct{\phi_{A}}{\phi_{B}}|=|t$|. (15)
We now derive the optimal minimum-error (Helstrom) discrimination strategy between $|\rho_{A}=\ket{\phi_{A}}\bra{\phi_{A}}$ and $|\rho_{B}=\ket{\phi_{B}}\bra{\phi_{B}}$, and we identify the optimal POVM explicitly.
Helstrom operator and minimum error probability, For priors $p_{1},p_{2}\geq 0$ with $p_{1}+p_{2}=1$, define the Helstrom operator
$\Delta:=p_{1}\rho_{A}-p_{2}\rho_{B}.$ (16)
The Helstrom theorem states that the minimum achievable average error probability is
$P_{err}^{\min}=\frac{1}{2}\Big(1-\|\Delta\|_{1}\Big),$ (17)
where $\|\cdot\|_{1}$ denotes the trace norm, and that the optimal two-outcome POVM is the projective measurement onto the positive and negative eigenspaces of $\Delta$, see more details Appendix C.
The eigenvalues of $\Delta$ are therefore
$\lambda_{\pm}=\pm\frac{1}{2}\sqrt{a^{2}+|b|^{2}}=\pm\frac{|r|}{2}.$ (18)
Since $\Delta$ is Hermitian, its trace norm is the sum of absolute eigenvalues,
$\|\Delta\|_{1}=|\lambda_{+}|+|\lambda_{-}|=|r|.$ (19)
Substituting Eq. (19) into Eq. (17) yields the Helstrom minimum error and maximum success probabilities:
$P_{err}^{\min}=\frac{1}{2}(1-|r|),\qquad P_{suc}^{\max}=1-P_{err}^{\min}=\frac{1}{2}(1+|r|).$ (20)
In terms of the overlap in Eq. (15) $s=|t|$ we have $|r|=\sqrt{1-|t|^{2}}=\sqrt{1-s^{2}}$, so
$P_{err}^{\min}=\frac{1}{2}\Big(1-\sqrt{1-s^{2}}\Big)=\frac{1}{2}\Big(1-\sqrt{1-|t|^{2}}\Big).$ (21)
It is natural to define the (equal-prior) Helstrom trace-distance distinguishability as (See more details in Appendix. D)
$D_{\mathrm{Hel}}:=\|\Delta\|_{1}=|r|=\sqrt{1-|t|^{2}}.$ (22)
Since the ZWM singles visibility is $V=|t|$, one obtains the saturated pure-marker duality relation
$D_{\mathrm{Hel}}^{2}+V^{2}=|r|^{2}+|t|^{2}=1.$ (23)
Thus, in the pure-marker regime, the distinguishability appearing in the complementarity relation coincides with the Helstrom distinguishability for minimum-error identification of the emitting crystal.
In the optimal Helstrom POVM (explicit eigenmodes). The optimal Helstrom POVM in Eq. (11) for equal priors is the projective measurement onto the eigenvectors of $\Delta$:
[…]
A convenient normalized eigenbasis of $\Delta$ is then
[…]
For completeness, for arbitrary priors $p_{1}\neq p_{2}$, the Helstrom bound for two pure states with overlap in Eq. (15) $|s=|\innerproduct{\phi_{A}}{\phi_{B}}$| is
$P_{err}^{\min}=\frac{1}{2}\Big(1-\sqrt{1-4p_{1}p_{2}\,s^{2}}\Big)=\frac{1}{2}\Big(1-\sqrt{1-4p_{1}p_{2}\,|t|^{2}}\Big).$ (26)
Equivalently,
$\|\Delta\|_{1}=\sqrt{1-4p_{1}p_{2}\,|t|^{2}}.$ (27)
For the equal-pump ZWM scenario $p_{1}=p_{2}=\tfrac{1}{2}$, these expressions reduce to Eqs. (20)–(22).
Equation (23) shows that, for pure idler marker states, wave–particle duality in the ZWM interferometer is nothing but the geometry of minimum-error quantum state discrimination. The visibility $V$ quantifies the wave character through the overlap of the two marker states. However, $D_{\mathrm{Hel}}$ quantifies the particle character as the optimal Helstrom distinguishability—i.e., the best possible ability to identify which crystal fired with minimum average error.
Importantly, this interpretation is operational: $D_{\mathrm{Hel}}$ is not an abstract distinguishability parameter but directly determines the optimal decision strategy. The error probability via the Helstrom bound (105). When the marker states are pure, the trade-off between interference visibility and which-crystal information is saturated. In the presence of loss or mixing (mixed marker states), the equality generally becomes an inequality, $D_{\mathrm{Hel}}^{2}+V^{2}\leq 1$, reflecting reduced operational distinguishability (See Appendix I).
Figure 2: Visibility $V(T)$, distinguishability $D(T)=\sqrt{1-V(T)^{2}}$, and the Helstrom minimum error probability $P_{\mathrm{err}}^{\min}(T)=\tfrac{1}{2}\big[1-D(T)\big]$ as functions of the object transmittance $T$ in the induced-coherence interferometer. Increasing transmittance enhances signal interference visibility while reducing which-crystal distinguishability encoded in the idler. The Helstrom bound quantifies the optimal minimum-error probability for inferring which nonlinear crystal emitted the photon pair from a single idler measurement. Illustrating the operational complementarity between interference and source discrimination.
In Fig. 2, the Helstrom error probability is transparent in the two extreme limits of the object transmittance. For an opaque object ($T=0$), the conditional idler states associated with emission from crystals A and B are orthogonal. If crystal A fires the idler is reflected, whereas if crystal B fires the idler bypasses the object. A single idler detection event therefore identifies the emitting crystal with certainty. The minimum error probability vanishes, $P_{\mathrm{err}}^{\min}=0$. Conversely, for a fully transparent object ($T=1$), both emission alternatives prepare the same idler state. The idler then carries no information about the source. Any measurement strategy reduces to random guessing, yielding $P_{\mathrm{err}}^{\min}=1/2$. Intermediate transmittance values interpolate smoothly between these limits ($0<T<1$), quantifying how partial which-crystal information degrades induced coherence.
V Zero-error strategy: IDP state discrimination
In this section, we treat the same “which-crystal fired?” problem, but now in the unambiguous (zero-error) discrimination setting [Barnett, Clarke2001IDP, Len2018IDP]: whenever the measurement makes a conclusive guess, it is never wrong, at the price of allowing an inconclusive outcome (See Appendix E).
In the Sec.II and Sec. III, two pure idler marker states are nonorthogonl, perfect identification is impossible. In the Idler measurement, any zero-error strategy must rely on unambiguous discrimination (USD) as shown in the Fig. 1. The optimal solution is the IDP measurement, whose minimal inconclusive probability is set by the state overlap.
The idler states $|\phi_{A}\rangle$ and $|\phi_{B}\rangle$ produced by the two crystals occupy only the single-photon idler subspace spanned by $|1,0\rangle$ and $|0,1\rangle$. In this two-dimensional subspace, each $|\phi_{j}\rangle$ admits a unique (up to phase) orthogonal complement $|\phi_{j}^{\perp}\rangle$.
We use the parametrization
$|\phi_{A}\rangle=r\,|1,0\rangle+t\,|0,1\rangle,\qquad|\phi_{B}\rangle=|0,1\rangle,$ (28)
with $|r|^{2}+|t|^{2}=1$.
The orthogonal complement of $|\phi_{B}\rangle$ is immediate:
$|\phi_{B}^{\perp}\rangle=|1,0\rangle.$ (29)
For $|\phi_{A}\rangle$, we write a general state in this subspace as $|\phi_{A}^{\perp}\rangle=\alpha\,|1,0\rangle+\beta\,|0,1\rangle$ and impose $\langle\phi_{A}^{\perp}|\phi_{A}\rangle=0$, giving
$\alpha^{*}r+\beta^{*}t=0.$ (30)
A convenient normalized solution is
$|\phi_{A}^{\perp}\rangle=t^{*}\,|1,0\rangle-r^{*}\,|0,1\rangle,$ (31)
where we used $|r|^{2}+|t|^{2}=1$.
An unambiguous discrimination measurement is a 3-outcome POVM [Barnett, Peres, Dieks, Clarke2001IDP]
[…]
with the decision rule:
[…]
  • outcome $\Pi_{inc}$: declare “Inconclusive (incinc)”.
Zero-error (unambiguous) constraints, unambiguous discrimination requires that a conclusive outcome never clicks for the wrong state:
[…]
Since $\rho_{j}$ are rank-1 projectors and $\Pi_{k}\geq 0$, Eq. (33) is equivalent to the support conditions
[…]
with parameters $a,b\geq 0$, and where
[…]
The inconclusive POVM element is then fixed by completeness:
[…]
The only remaining requirement is positivity:
[…]
The inconclusive probabilities are
$\begin{split}Q_{A}\equiv&P(inc|A)=1-P(A|A),\\ Q_{B}\equiv&P(inc|B)=1-P(B|B),\end{split}$ (41)
and the average inconclusive probability is
$Q\equiv P(inc)=p_{1}Q_{A}+p_{2}Q_{B}.$ (42)
The IDP problem is: minimize $Q$ subject to $\Pi_{inc}\geq 0$.
In the Schrödinger-picture derivation see more details Appendix F. Unambiguous discrimination between $|\phi_{A}\rangle$ and $|\phi_{B}\rangle$ with equal priors requires a three-outcome POVM $\{\Pi_{A},\Pi_{B},\Pi_{inc}\}$. The corresponding POVM, to conclusive identification of $|\phi_{A}\rangle$, conclusive identification of $|\phi_{B}\rangle$, and an inconclusive result. Physically, such a POVM can be realized by embedding the two-dimensional idler subspace spanned by $\{|1,0\rangle_{I_{A}^{{}^{\prime}},I_{B}},|0,1\rangle_{I_{A}^{{}^{\prime}},I_{B}}\}$ into a three-dimensional single-photon space. By introducing a third auxiliary mode $C$ initially in the vacuum state $|0_{C}\rangle$.
In the single-photon sector of the three modes $(I_{A}^{{}^{\prime}},I_{B},C)$ we use the basis
[…]
$|0,1,0\rangle,$
$|0,0,1\rangle\}_{I_{A}^{{}^{\prime}},I_{B},C}.$
  • $p_{1}$ is the probability that input $|\phi_{A}\rangle$ produces a click in mode 1 (conclusive “crystal A”),
  • $p_{2}$ is the probability that input $|\phi_{B}\rangle$ produces a click in mode 2 (conclusive “crystal B”),
The form of Eqs. (146)–(147) enforces the zero-error conditions: for $|\phi_{A}\rangle$ there is no amplitude in $|010\rangle$ (the “B” detector never fires), and for $|\phi_{B}\rangle$ there is no amplitude in $|100\rangle$ (the “A” detector never fires).
The average success probability of the IDP measurement is then
$P_{suc}^{\mathrm{opt}}=\tfrac{1}{2}p_{1}+\tfrac{1}{2}p_{2}=1-s,$ (46)
and the optimal inconclusive probability is
$P_{inc}^{\mathrm{opt}}=1-P_{suc}^{\mathrm{opt}}=s=|\langle\phi_{A}|\phi_{B}\rangle|.$ (47)
In the ZWM interferometer, $s=|\langle\phi_{A}|\phi_{B}\rangle|=|t|=V$, so the IDP theory yields
$P_{inc}^{\mathrm{opt}}=V,\qquad P_{suc}^{\mathrm{opt}}=1-V.$ (48)
$D^{2}+\bigl(P_{inc}^{\mathrm{opt}}\bigr)^{2}=1$ (49)
for pure marker states, and establishes the interpretation $V=P_{inc}^{\mathrm{opt}}$ used in the main text.
Figure 3: (a) Operational complementarity between visibility $V$, distinguishability $D$, and the optimal IDP failure probability $P_{inc}^{\mathrm{opt}}$ for pure marker states with overlap $s=|\langle\phi_{A}|\phi_{B}\rangle|$ (see Appendix F). The curves satisfy $D^{2}+V^{2}=1$ and $V=P_{inc}^{\mathrm{opt}}$, identifying wave–particle duality with the optimality condition of unambiguous state discrimination. (b) Comparison between distinguishability obtained via the coincidence-based method of Torres et al. [Torres2024] (solid) and via optimal IDP discrimination on the idler (dashed). The coincidence method saturates at $D_{\mathrm{coinc}}(T=1)\approx 0.52$, whereas IDP extracts greater distinguishability for intermediate $0<T<1$. Both approaches correctly give $D=0$ at $T=1$, where the marker states become identical. The advantage of IDP therefore appears only for partially overlapping markers.
Figure 3 provides a direct operational comparison between distinguishability extracted from signal-arm visibility (Torres et al.) [Torres2024] and distinguishability obtained via optimal IDP discrimination performed on the idler. In panel (a), we plot the complementarity relation for pure which-crystal markers, showing that the visibility $\mathcal{V}$, the optimal IDP failure probability $P_{inc}^{\mathrm{opt}}$. The particle-like distinguishability $D$ are constrained by the operational identity $\mathcal{V}=P_{inc}^{\mathrm{opt}}$ and $D^{2}+\mathcal{V}^{2}=1$. This demonstrates that wave–particle duality is exactly equivalent to an optimal unambiguous state discrimination task for the which-path markers.
Panel (b) compares distinguishability extracted from the coincidence-based method of Torres et al. [Torres2024] (solid curve). The distinguishability achievable via optimal Ivanovic–Dieks–Peres (IDP) unambiguous discrimination on the idler (dashed curve), both expressed in the same metric. In the Torres approach, $D$ is inferred from the imbalance between second-order correlations $g^{(2)}_{13}$ and $g^{(2)}_{23}$. It saturates at a finite value $D_{\mathrm{coinc}}(T=1)<1$ even when the idler transmissivity $T$ is unity. In the opposite limit, the IDP strategy operates directly on the which-crystal idler marker states and achieves the optimal extraction of available which-path information. Both methods converge to $D=0$ when $T=1$, since the idler marker states become identical and no measurement—IDP or otherwise—can distinguish the emission events. The advantage of IDP emerges in the intermediate regime $0<T<1$, where coincidence-based extraction underestimates the distinguishability available in principle. This highlights that path information in induced coherence is fundamentally limited by measurement strategy rather than interferometer geometry.
VI General Marker state
@@
The same reasoning applies to any balanced two-path interferometer in the single-excitation regime with pure which-path markers $|\mu_{A}\rangle$ and $|\mu_{B}\rangle$. The fringe visibility is determined by their overlap (See Appendix G),
$V=|\langle\mu_{A}|\mu_{B}\rangle|,$
@@
$D=\sqrt{1-|\langle\mu_{A}|\mu_{B}\rangle|^{2}}.$
For equal priors, the optimal IDP strategy yields $P_{inc}^{\mathrm{opt}}=|\langle\mu_{A}|\mu_{B}\rangle|$, so that
$D^{2}+\big(P_{inc}^{\mathrm{opt}}\big)^{2}=1$
@@
Table 1: Operational meaning of visibility and distinguishability as quantum decision tasks.
Strategy Outcomes Errors Operational quantity What $V$ means
Previous ZWM setup 11 No errors $D=\sqrt{1-V^{2}}$ $V$ (signal measurement)
Helstrom (min-error) 22 Allowed (minimized) $D_{\mathrm{Hel}}=\tfrac{1}{2}\|\rho_{A}-\rho_{B}\|_{1}$ $V=|\langle\phi_{A}|\phi_{B}\rangle|$ (pure markers)
USD / IDP (zero-error) 33 None (if conclusive) $D_{\mathrm{IDP}}=\sqrt{1-\bigl(P_{\mathrm{inc}}^{\mathrm{opt}}\bigr)^{2}}$ $V=P_{\mathrm{inc}}^{\mathrm{opt}}$ (pure markers, equal priors)
Mixed markers 22 or 33 Unavoidable Reduced ($D<1$) $V\leq F(\rho_{A},\rho_{B})$
VII Experimental Protocol
Schematic of the ZWM induced-coherence interferometer as shown in Fig. 1, in the idler measurement implementing unambiguous state discrimination. The idler modes are routed into an interferometric network of beam splitters (BS) and phase shifters (PS), realizing the optimal Ivanovic–Dieks–Peres (IDP) three-outcome POVM [Bennett1996, Len2018IDP, Len2018IDP]. Detection events at $D_{1}$ ($D_{2}$) unambiguously identify crystal A (B) and $D_{inc}$ corresponds to an inconclusive outcome. No ancillary modes beyond vacuum inputs are required.